OpenStudy (anonymous):

perfect square trinomials x^2 + 6x +36

OpenStudy (xapproachesinfinity):

what do you need to do? factor it?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

do you know this (x+a)(x+b)=x^2+(a+b)x+ab ca you find what the two number if you multiply them you get 36 and if you add them you get 6

OpenStudy (anonymous):

thank you very much till next time :)

OpenStudy (anonymous):

x^2+bx+c , where c = (b/2)^2 is a perfect square.

OpenStudy (xapproachesinfinity):

It doesn't seem a perfect square by the way did you miss something

OpenStudy (xapproachesinfinity):

Check if you wrote properly, because it is not a perfect square

OpenStudy (anonymous):

That's what I thought

OpenStudy (xapproachesinfinity):

Lol I just realized when I explained to her the method to factor any trinomial that is factorable

OpenStudy (anonymous):

lets try another 9x^2 + 24x - 16

OpenStudy (xapproachesinfinity):

please post the complete question. If a question is not complete we won't know what we want to do and how to go with it! understanding the question is key factor to answering, it's half the answer